what is the ordered pair of 2y-10x=-46 and x-3y=13
Answer by checkley77(12844)

2y-10x=-46
x-3y=13 MULTIPLY THIS EQUATION BY 10 & ADD.
10X-30Y=130
-10X+2Y=-46
--------------
-28Y=84
Y=84/-28
Y=-3 ANS.
2*-3-10X=-46
-6-10X=-46
-10X=-46+6
-10X=-40
X=-40/-10
X=4 ANS.
PROOF:
4-3*3=13
4+9=13
13=13